You are considering how to invest part of your retirement savings.You have decided to put $400,000

into three​ stocks 59% of the money in GoldFinger​ (currently $ 25​/share), 17 % of the money in Moosehead​ (currently

$ 97/share),and the remainder in Venture Associates​ (currently $ 10​/share).

Suppose GoldFinger stock goes up to $34​/share,

Moosehead stock drops to $60​/share, and Venture Associates stock rises to $ 19 per share.

a. What is the new value of the​ portfolio?

b. What return did the portfolio​ earn?

c. If you​ don't buy or sell any shares after the price​ change, what are your new portfolio​ weights?
Answer #1
Shares inv. in Goldfinger = Amnt inv.*%age in Goldfinger/price of Goldfinger=400000*0.59/25=9440
Shares inv. in Moosehead = Amnt inv.*%age in Moosehead/price of Moosehead=400000*0.17/97=701.0309
Shares inv. in Venture ass. = Amnt inv.*%age in Venture ass./price of Venture ass.=400000*0.24/10=9600
A.
New portfolio value= Goldfinger new price*Goldfinger shares + Moosehead new price*Moosehead shares + Venture ass. new price*Venture ass. shares
=34*9440+60*701.0309+19*9600
=545421.85
B.
Return % = New portfolio value/old portfolio value-1
=545421.85/400000-1
=36.36%
C.
New weight of Goldfinger = New price*shares/new porfolio value
=34*9440/545421.85
=0.5885
New weight of Moosehead = New price*shares/new porfolio value
=60*701.0309/545421.85
=0.0771
New weight of Venture ass. = New price*shares/new porfolio value
=19*9600/545421.85
=0.3344